Classify the financial problem. Assume a 11% interest rate compounded annually.
Find the value of a $1,000 certificate in 6 years. (Round your answer to the nearest cent.)

Answer by lwsshak3(11628)

Classify the financial problem. Assume a 11% interest rate compounded annually.
Find the value of a $1,000 certificate in 6 years. (Round your answer to the nearest cent.)
***
Compound Interest Formula: A=P(1+i)^n, P=initial investment, i=interest rate per period, A=amt after n-periods.
For given problem:
P=1000
i=0.11
n=6
..
A=1000(1+0.11)^6
A=1000(1.11)^6
A=1000*1.87042 (rounded)
A=1870.42
value of a $1,000 certificate in 6 years at 11% interest rate compounded annually=$1872.42